OpenStudy (howard-wolowitz):

Find the sum of the geometric sum: 12+36+108+ ... + 78,732

OpenStudy (howard-wolowitz):

354,288 118,092 39,360 1,062,876

OpenStudy (howard-wolowitz):

@freckles

OpenStudy (howard-wolowitz):

hi can you help me

OpenStudy (welshfella):

sum of n terms = a1 * r^n - 1 ------ r - 1 r = common ratio = 36/12 = 3, a1 = first term = 12

OpenStudy (howard-wolowitz):

ok

OpenStudy (welshfella):

now we need to find the value of n the general form is a1 , a1*r , a*1r^2

OpenStudy (howard-wolowitz):

ok

OpenStudy (welshfella):

so the nth term = a1 r^(n-1)

OpenStudy (howard-wolowitz):

you know i actually like you u actually show me hoe and not except me to do it

OpenStudy (howard-wolowitz):

you actually show how to do it so for that I THANKS YOU

OpenStudy (howard-wolowitz):

most poeple just say the formula crap

OpenStudy (welshfella):

well let 78,732 = 12* r^(n-1) we need to solve this for n

OpenStudy (howard-wolowitz):

ok whats r

OpenStudy (welshfella):

oh sorry r = 3

OpenStudy (howard-wolowitz):

ok

OpenStudy (howard-wolowitz):

then what do i do

OpenStudy (welshfella):

Ok divide both sides by 12 gives 3^(n-1) = 6561

OpenStudy (welshfella):

now you can find the value of n by taking logs of both sides can you do that?

OpenStudy (howard-wolowitz):

yes i got the A option

OpenStudy (howard-wolowitz):

also just know that mathman said it was C

OpenStudy (howard-wolowitz):

so what did you get dude

OpenStudy (welshfella):

I dont think its either taking logs (n -1) ln 3 = ln 6561 n - 1 = ln6561 / ln 3 = 8 n = 9

OpenStudy (welshfella):

so S9 = 12 * 3^9 - 1 ------ 3 - 1

OpenStudy (welshfella):

can you work that out?

OpenStudy (welshfella):

3^9 = 19683 so subtract 1 from that then divide by 2 then multiply by 12 and you have the answer

OpenStudy (howard-wolowitz):

yes one sec

OpenStudy (howard-wolowitz):

you cant do that its a deciaml

OpenStudy (welshfella):

??

OpenStudy (howard-wolowitz):

u cant do that

OpenStudy (howard-wolowitz):

u get a decimal

OpenStudy (howard-wolowitz):

dude just tell me what you got

OpenStudy (howard-wolowitz):

@freckles

OpenStudy (howard-wolowitz):

i keep getting decimals

OpenStudy (welshfella):

I can tell you the answer but I'd rather you work it out from the expression I gave you its 12 * (3^9 - 1) / 2 just grab ur calculator and compute it

OpenStudy (freckles):

\[12 \cdot \frac{1-3^9}{1-3}\] what did you get when you input into the calculator @Howard-Wolowitz ?

OpenStudy (freckles):

this one actually comes out to be a pretty whole number

OpenStudy (howard-wolowitz):

let me do that

OpenStudy (welshfella):

yes

OpenStudy (howard-wolowitz):

i got -19685

OpenStudy (welshfella):

no are you doing the calculations in the right order/ first work out 3^9 then subtract 1 from the result then divide by 2 finally multiply by 12

OpenStudy (freckles):

what kinda calculator are you using? it is a ti-83 or one of those one entry type calculators ?

OpenStudy (howard-wolowitz):

my phone

OpenStudy (freckles):

ewww so it is most likely one entry can your phone do exponents?

OpenStudy (howard-wolowitz):

ok i did the frist step and got 19683

OpenStudy (howard-wolowitz):

then the second step got me to 19682

OpenStudy (freckles):

so then you ahve this \[12 \cdot \frac{1-19683}{1-3}\]

OpenStudy (welshfella):

thats correct now subtarct 1 from that

OpenStudy (freckles):

\[12 \cdot \frac{-19682}{1-3}\]

OpenStudy (welshfella):

or you can do its freckles way

OpenStudy (freckles):

either way it doesn't matter

OpenStudy (howard-wolowitz):

the third i got a decimal

OpenStudy (welshfella):

it will give the same result

OpenStudy (freckles):

wait how did you get a decimal ?

OpenStudy (howard-wolowitz):

i cant do it that way freckles on my phone

OpenStudy (freckles):

1-3=?

OpenStudy (howard-wolowitz):

-2

OpenStudy (freckles):

\[12 \cdot \frac{-19682}{-2} \\ 12 \cdot \frac{19682}{2}\]

OpenStudy (freckles):

so how did you get a decimal here ?

OpenStudy (howard-wolowitz):

third is 9841 1/2

OpenStudy (freckles):

how did you get that

OpenStudy (howard-wolowitz):

or 19683/2

OpenStudy (freckles):

\[12 \cdot \frac{19682}{2} \text{ is the same as } \frac{12}{2} \cdot 19682 \] without using a calculator tell me what is 12/2 ?

OpenStudy (welshfella):

- you dont find a third - you divide by 2 the 3 -1 on bottom is worked out first before you do the divide

OpenStudy (howard-wolowitz):

ok so now just do that ok one sec

OpenStudy (freckles):

\[12 \cdot \frac{1-3^9}{1-3}=\frac{12}{1-3} \cdot (1-3^9)=\frac{12}{-2}(-19682) \\ =\frac{12}{2}(19682)=6(19682)=?\]

OpenStudy (howard-wolowitz):

Ok I got B

OpenStudy (howard-wolowitz):

118,092

OpenStudy (howard-wolowitz):

is that right

OpenStudy (freckles):

ok and you might want to practice on the actual calculator that you can use on a test

OpenStudy (freckles):

pretty sure most instructors don't let you use phones on test